OpenStudy (anonymous):

Pre calc question~ Help please! So, for the question below, I'm supposed to mark any solutions for the equation: tan^2x-3=0. My options are: A. 2pi/3 B. pi C. -pi/3 D. 5pi/3 I keep coming up with A and C as the answer, however those choices aren't correct.

OpenStudy (anonymous):

@ganeshie8 help please? :)

OpenStudy (nincompoop):

like this? \[\tan^{2x-3}=0\]

OpenStudy (anonymous):

@nincompoop Here's what it looks like on my assignment!

hero (hero):

One way to do it is to check each one using this approach: tan^2(2pi/3)-3=0 tan^2(pi)=0 tan^2(-pi/3)-3=0 tan^2(5pi/3)-3=0

hero (hero):

The other way to do it is to actually solve for x: \(\tan(x)^2 - 3 = 0\) \(\tan(x)^2 = 3\) \(\tan(x) = \pm\sqrt{3}\) \(x = \tan^{-1}(\pm \sqrt{3})\)

OpenStudy (anonymous):

For the first approach, I'd just plug the solutions into the equation? and if they come out whole, they'd be an answer?

hero (hero):

For the first approach, a given expression must equal zero to be correct.

hero (hero):

Be sure to put your calc in radian mode

OpenStudy (anonymous):

Got it! That makes sense now. Thank you :)

OpenStudy (anonymous):

Wait, so now all I have for solutions is pi?

hero (hero):

I was going to mention that if you solved it the other way you must add the period of tan to your original solutions to get other values

OpenStudy (anonymous):

Oh. Sorry, I'm very bad at math in this form, how would I add the period of tan into the whole thing?

OpenStudy (anonymous):

Ah, okay. When I only used pi as an answer, it said I needed one more... Which is why I'm so confused! Thank you for your help though, at least I know how to approach similar problems now.

hero (hero):

What's interesting is you mentioned that you originally got A and C as the correct choices. That's the same thing I get when I check.

hero (hero):

But I also get one more as well

OpenStudy (anonymous):

Oh! Did you get D as the third one?

hero (hero):

Yes and desmos.com confirms: https://www.desmos.com/calculator/a898xwnb3s

OpenStudy (anonymous):

Thank you so much!

hero (hero):

Just to emphasize, I plotted some points on the graph here: https://www.desmos.com/calculator/5uq6uammlf

OpenStudy (anonymous):

The visual does help me understand it more.

OpenStudy (anonymous):

:)